### v2

# Counting rational points on smooth quartic and quintic surfaces

Abstract.
Let $X\subseteq\mathbb{P}^{3}$ be a smooth projective surface of degree $d\geq 4$ defined over a number field $K$, and let $N_{X^{\prime}}(B)$ be the number of rational points of $X$ of height at most $B$ that do not lie on lines contained in $X$. Assuming a suitable hypothesis on the size of the rank of Abelian varieties, we show that $N_{X^{\prime}}(B)\ll_{K,d,\varepsilon}B^{4/3+\varepsilon}$ for any fixed $\varepsilon>0$. This improves an unconditional bound from Salberger for $d=4$ and $d=5$. The proof, based on an argument of Heath-Brown, consists of cutting $X$ by projective planes and using a uniform version of Faltings’s Theorem, due to Dimitrov, Gao, and Habegger, to bound the number of rational points on the plane sections of $X$. More generally, we prove that if $X\subseteq\mathbb{P}^{n}$ is a non-degenerate non-uniruled smooth projective surface defined over $K$, then $N_{X^{\prime}}(B)\ll_{K,n,d,\varepsilon}B^{\frac{n+1}{n}+\varepsilon}$.
1. Introduction
Let $K$ be a number field, and let $H_{K}$ be the height function in $\mathbb{P}^{n}$ given by:
$H_{K}([x_{0}\colon\cdots\colon x_{n}])=\prod_{v}\max\{\lVert x_{0}\rVert_{v},\dotsc\lVert x_{n}\rVert_{v}\},$
where $v$ runs through the places of $K$ and $\lVert x\rVert_{v}$ is the normalized absolute value associated with the place $v$ (this is the height function defined in [10, Section B.2]).
We consider a smooth hypersurface $X\subseteq\mathbb{P}^{n}$ of degree $d$, and we want to bound the number of $K$-rational points of bounded height on $X$. For this, define the counting function
$N_{X}(B)\coloneqq\#\{x\in X(K)\colon H_{K}(x)\leq B\}.$
For $K=\operatorname{\mathbb{Q}}$, Verzobio shows in [21] that for $n\geq 4$ and $d\geq 50$ we have
$N_{X}(B)\ll_{n,d,\varepsilon}B^{n-2+\varepsilon},$
and gives other bounds for $n\geq 4$ and $d\geq 6$. It is conjectured that this bound holds for all $n\geq 4$ and $d\geq 3$.
For $n=3$, it is clear that this bound does not hold if $X$ contains a line defined over $K$, since in this case we would have $N_{X}(B)\gg B^{2}$. Therefore, when $X\subseteq\mathbb{P}^{3}$ is a smooth surface, we are interested in counting the rational points on the set $X^{\prime}\subseteq X$, where $X^{\prime}$ is the complement of the union of all lines on $X$. By Lemma 2.3, for $d\geq 4$, $X^{\prime}$ is obtained from $X$ by removing a finite number of lines. We define the counting function
$N_{X^{\prime}}(B)\coloneqq\#\{x\in X^{\prime}(K)\colon H_{K}(x)\leq B\}.$
For $n=3$ and $K=\operatorname{\mathbb{Q}}$, Salberger shows in [18, Theorem 0.5] the bound
$N_{X^{\prime}}(B)\ll_{d}B^{3/\sqrt{d}}(\log B)^{4}+B.$ (1)
We aim to improve this bound for $d=4$ and $5$, using a technique developed by Heath-Brown in [8] and assuming a conjectural bound for the rank of Abelian varieties. In his paper, Heath-Brown shows that for $d=3$ and assuming such a hypothesis, we have, for any fixed $\varepsilon>0$, a bound
$N_{X^{\prime}}(B)\ll_{X,\varepsilon}B^{4/3+\varepsilon}.$
This technique was also applied by Glas and Hochfilzer to bound the number of rational points on Del Pezzo surfaces of low degree over a global field, assuming the validity of the same conjectural bound for number fields ([6]).
We will cover the points of bounded height with projective planes of bounded height in the dual projective space $(\mathbb{P}^{3})^{\vee}$. This can be done using a result from Schmidt ([19]). We then apply another result from the same paper to bound the number of such planes asymptotically, and proceed by estimating the number of points of bounded height in the intersection between $X$ and each of these planes. This intersection will be generically a smooth curve. In this case, we use a uniform version of Faltings’s Theorem, due to Dimitrov, Gao, and Habegger:
Lemma 1.1 (Uniform Faltings’s Theorem ([5])).
For each integer $g\geq 2$, there exists a constant $c(g)>0$ depending on $K\colon\operatorname{\mathbb{Q}}$] such that, for every smooth projective curve $C$ over $K$ of (geometric) genus $g$, we have
${C(K)}\leq c(g)^{1+\rho(C)},$
where $\rho(C)$ denotes the rank of $\operatorname{\operatorname{Jac}}(C)(K)$.
This is [5, Theorem 1]. Compare it also with [12, Theorem 4], which eliminates the dependence on the degree $K:\operatorname{\mathbb{Q}}]$.
We bound the number of rational points on curves of genus at least $2$ that appear as irreducible components of the intersections between $X$ and rational planes by applying the above result to their normalizations.
We can proceed similarly to bound the number of rational points on the irreducible components of genus $1$, by using a bound on the number of smooth cubic plane curves due to Heath-Brown ([8]), which is generalized to global fields by Glas and Hochfilzer ([6]). Finally, we show that the curves of genus 0 that appear do not alter our estimation. We remove dependence on $X$ by applying a standard trick due to Heath-Brown.
To reach our goal, we require the following hypothesis:
Hypothesis 1.2 (Rank Hypothesis).
For any $K$-Abelian variety $A$, let $N_{A}$ denote its conductor (that is, the ideal norm of the conductor ideal of $A$) and let $r_{A}$ denote its rank. Then we have $r_{A}=o(\log N_{A})$ as $N_{A}\to\infty$.
This hypothesis is formulated in [8] for elliptic curves over the rational numbers.
We remark that what we actually need is the a priori weaker bound $r_{A}=o(h_{\operatorname{Falt}}(A))$, where $h_{\operatorname{Falt}}(A)$ denotes the Faltings’s height of $A$. However, the plausibility of the Rank Hypothesis is suggested by the following analytical argument:
Assuming modularity for the Abelian variety $A$ over the number field $K$, this hypothesis would hold for $A$ by a combination of the Generalized Riemann Hypothesis and the Birch and Swinnerton-Dyer conjecture. Indeed, let $L(A,s)$ be the $L$-function associated with the Abelian variety $A$. Then, if we define
$\Lambda(A,s)\coloneqq N_{A}^{s/2}((2\pi)^{-s}\cdot\Gamma(s))^{\dim A}\cdot L(A,s),$
by the modularity of $A$ the function $\Lambda(A,s)$ extends to an entire function and satisfies the functional equation
$\Lambda(A,2-s)=\pm\Lambda(A,s).$
Assuming the Generalized Riemann Hypothesis and proceeding by classical analytic arguments, this functional equation would give us
$\operatorname{ord}_{s=1}(L(A,s))=O\left(\frac{\log N_{A}}{\log\log N_{A}}\right)=o(\log N_{A}).$
Assuming the Birch and Swinnerton-Dyer conjecture, we obtain
$\operatorname{ord}_{s=1}(L(A,s))=r_{A},$
which gives the Rank Hypothesis. See [13, Proposition II.1] for a detailed proof of these implications in the case of the $L$-function of a modular form.
We now cite some known results that go in the direction of the Rank Hypothesis. By [14, Theorem 1], for an Abelian variety $A/K$ we have, unconditionally,
$r_{A}=O_{K,\dim A}(N_{A}).$
In the case of elliptic curves, more evidence is known. Notice that all elliptic curves over the rational number are modular by [1]. If we assume that an elliptic curve $E/\operatorname{\mathbb{Q}}$ has at least one rational point of order $2$, then it is known that (see [8, Section 6])
$r_{E}=O\left(\frac{\log N_{E}}{\log\log N_{E}}\right)=o(\log N_{E}).$
Our main result is the following:
Theorem 1.3.
Assume that the Rank Hypothesis holds. Let $X\subseteq\mathbb{P}^{3}$ be a smooth surface of degree $d\geq 4$, and let $\varepsilon>0$. Then we have $N_{X^{\prime}}(B)\ll_{K,d,\varepsilon}B^{4/3+\varepsilon}$.
This exponent is smaller than the one in (1) for $d=4$ and $d=5$, thereby improving the known bound in these cases.
Another advantage of our method is that it works over any number field, so we obtain the same bound for every choice of $K$. To do this, we use results of Paredes and Sasyk ([15]) and Glas and Hochfilzer ([6]). However, notice that we use the height $H_{K}$ of the field $K$, while in [15] they use the absolute height $H=H_{K}^{1/[K:\operatorname{\mathbb{Q}}]}$.
We can generalize our result for smooth surfaces contained in some bigger projective space:
Theorem 1.4.
Assume that the Rank Hypothesis holds. Let $n\geq 3$, let $X\subseteq\mathbb{P}^{n}$ be a non-degenerate non-uniruled smooth surface of degree $d$, and let $\varepsilon>0$. Then we have $N_{X^{\prime}}(B)\ll_{K,n,d,\varepsilon}B^{\frac{n+1}{n}+\varepsilon}$.
The proof of this result is analogous to the proof of Theorem 1.3, except that we cover $X$ by hyperplanes of $\mathbb{P}^{n}$ instead of by planes of $\mathbb{P}^{3}$. Since a uniruled surface always has Kodaira dimension -$\infty$, the above theorem applies to surfaces of general type, general elliptic surfaces, abelian surfaces, hyperelliptic surfaces, Enriques surfaces and K3 surfaces.
We hope to generalize this method in a future paper to obtain results about higher dimensional varieties.
Some articles use a slightly different definition of $H_{K}$, taking the Archimedean norm in infinite places instead of the maximum norm. However, since both norms differ by a multiplicative constant, this distinction will not matter in what follows, so we shall ignore it.
Given a polynomial $P(x_{0},\dotsc,x_{n})\in K[x_{0},\dotsc,x_{n}]$, we define its height $H_{K}(P)$ as the height over $K$ of the projective point whose entries are the coefficients of $P$, and its logarithmic height to be $h_{K}(P)\coloneqq\log H_{K}(P)$.
By genus, we always mean the geometric genus, unless explicitly stated. Besides, we will denote by $x_{0},x_{1},x_{2},x_{3}$ the coordinates of $\mathbb{P}^{3}$ and by $F(x_{0},x_{1},x_{2},x_{3})$ the form of degree $d$ that defines $X$.
Following [6], we define, for $x=(x_{0},\dotsc,x_{n})\in\mathcal{O}_{K}^{n+1}$:
$\lVert x\rVert\coloneqq\max_{v\mid\infty}\max\{\lVert x_{0}\rVert_{v},\dotsc,\lVert x_{n}\rVert_{v}\},$
and
$\lVert x\rVert_{\infty}\coloneqq\prod_{v\mid\infty}\max\{\lVert x_{0}\rVert_{v},\dotsc,\lVert x_{n}\rVert_{v}\}.$
Fix $\mathfrak{a}_{1},\dotsc,\mathfrak{a}_{h}\subseteq\mathcal{O}_{K}$ to be a full set of representatives in the class group of $K$. Whenever we indicate that an implied constant depends on $K$, we implicitly allow it to also depend on our choice of the ideal class group representatives. Define
$Z_{n}^{\prime}\coloneqq\{x\in\mathcal{O}_{K}^{n+1}\setminus\{0\}\colon(x_{0},\dotsc,x_{n+1})=\mathfrak{a}_{i}\text{ for some }i=1,\dotsc,h\}.$
Let $s_{K}$ be the number of infinite places of $K$. By [6, Lemma 2.8], there exist constants $c_{1},c_{2}>0$ depending on $K$ such that every member of $\mathbb{P}^{n}(K)$ has a representative $x\in Z_{n}^{\prime}$ such that
$c_{1}\lVert x\rVert\leq\lVert x\rVert_{\infty}^{1/s_{K}}\leq c_{2}\lVert x\rVert.$
Hence, if we define
$Z_{n}\coloneqq\{x\in Z_{n}^{\prime}\colon c_{1}\lVert x\rVert\leq\lVert x\rVert_{\infty}^{1/s_{K}}\leq c_{2}\lVert x\rVert\},$
then every point of $\mathbb{P}^{n}(K)$ has a representative in $Z_{n}$. Notice that, for $x\in Z_{n}^{\prime}$, $H_{K}(x)$ is of the same order as $\lVert x\rVert_{\infty}$, since $H_{K}(x)$ is equal to $\lVert x\rVert_{\infty}$ divided by the norm of the ideal $(x_{0},\dotsc,x_{n+1})=\mathfrak{a}_{i}$ for some $1\leq i\leq h$.
2. Proof of Theorem 1.3
We start with the following dichotomy:
Lemma 2.1.
Either $H_{K}(F)\ll_{K}B^{d\binom{d+3}{3}}$ or there exists a surface $Y\subseteq\mathbb{P}^{3}$ of degree $d$ such that $X\cap Y$ is a curve, and all $K$-rational points of $X$ of height smaller than $B$ are contained in this curve.
Proof.
This is a generalization of [3, Lemma 5] to number fields. We can assume that $F$, the form defining $X$, has coefficients in $Z_{\binom{d+3}{3}-1}$. Suppose that $x^{(1)},\dotsc,x^{(N)}$ are all the rational points of $X$ satisfying $H_{K}(x)\leq B$, and write these points with representatives in $Z_{3}$. Let $M$ be the $\binom{d+3}{3}\times N$ matrix whose $i$-th row consists of the $\binom{d+3}{3}$ possible monomials of degree $d$ in the variables $x_{0}^{(i)},\dotsc,x_{3}^{(i)}$.
If $f$ is the vector given by the coefficients of $F$, then $Mf=0$, which shows that the rank of $M$ is at most $\binom{d+3}{3}-1$. Thus, we can find a nonzero vector $g\in Z_{\binom{d+3}{3}-1}$ constructed from subdeterminants of $M$ such that $Mg=0$. Hence $\lVert g\rVert\ll B^{d\binom{d+3}{d}/s_{K}}$. Let $G$ be the form of degree $d$ corresponding to the vector $g$, and $Y\subseteq\mathbb{P}^{3}$ be the surface defined by $G$. Then $G(x^{(i)})=0$ for all $1\leq i\leq N$, hence $\{x^{(1)},\dotsc,x^{(N)}\}\subseteq X\cap Y$. If $G$ is not proportional to $F$, we are done. Otherwise, since both $f$ and $g$ are in $Z_{\binom{d+3}{3}-1}$, we have
$\lVert f\rVert\ll\lVert g\rVert\ll B^{d\binom{d+3}{3}/s_{K}}.$
Finally, we conclude that
$H_{K}(F)=H_{K}(f)\ll_{K}\lVert f\rVert_{\infty}\ll_{K}\lVert f\rVert^{s_{K}}\ll B^{d\binom{d+3}{3}}$.
If we are in the second case of the lemma above, then $N_{X^{\prime}}(B)\leq N_{(X\cap Y)^{\prime}}(B)$, where $X\cap Y$ is a curve. This curve has a finite number of components. Let $C\subseteq X\cap Y$ be an irreducible component of degree $e\geq 2$. By [15, Theorem 1.8],
$N_{C}(B)\ll_{K,d,\varepsilon}B^{2/e+\varepsilon}\leq B^{1+\varepsilon}.$
If $e=1$, then $C$ is a line, and by definition it does not contribute to the calculation of $N_{X^{\prime}}(B)$. In this case, we conclude that $N_{X^{\prime}}\ll_{K,d,\varepsilon}B^{1+\varepsilon}<B^{4/3+\varepsilon}$. Therefore, from now on, we will assume that we are in the first case of the lemma above, that is, $H_{K}(F)\ll B^{d\binom{d+3}{3}}$.
Let $V$ be an $n$-dimensional $K$-vector space, and let $v\in V$. Let $H_{K}$ be the height on $V$ given by the identification $V=\operatorname{\mathbb{A}}_{K}^{n}(K)\xhookrightarrow{}\mathbb{P}^{n}(K)$. Let $1\leq k\leq n-1$. We start by defining the height in the Grassmannian
$\operatorname{\operatorname{Gr}}_{K}(k,n)\coloneqq\{W\subseteq V\colon W\text{ is a $K$-linear subspace of dimension }k\}.$
In our argument, we will only use $\operatorname{\operatorname{Gr}}_{K}(3,4)$, which is the dual projective space $(\mathbb{P}^{3})^{\vee}$. However, the more general concept of a Grassmanian will be useful if one wants to generalize this argument for varieties of higher dimension.
Let $W\subseteq V$ be a $k$-dimensional subspace. Let $w_{1},\dotsc,w_{k}$ be any basis of $W$. Consider the exterior product $w_{1}\wedge\cdots\wedge w_{k}\in\bigwedge^{k}V\cong K^{\binom{n}{k}}$. We define $H_{K}(W)\coloneqq H_{K}(w_{1}\wedge\cdots\wedge w_{k})$. It is possible to show that this height does not depend on the choice of basis, so $H_{K}(W)$ is well-defined. For more on the heights of subspaces, see [19].
Returning to our original problem, let $x\in X(K)$. We want to find a projective plane $\Pi\cong\mathbb{P}^{2}$ of small height containing $x$. This is equivalent to finding a $3$-dimensional subspace of a $4$-dimensional space that contains a given vector $x$. By [19, Theorem 2], there exists such a plane $\Pi$ with $H_{K}(\Pi)\ll_{K}B^{1/3}$. This shows that, for a given $B>0$, the set
$\{x\in X(K)\colon H_{K}(x)\leq B\}.$
can be covered by planes of height $O_{K}(B^{1/3})$. Using now [19, Theorem 3] (or the more precise version [20, Theorem 1]), we see that the number of such planes is
$O_{K}\big(\big(B^{1/3}\big)^{4}\big)=O_{K}(B^{4/3}).$
We proceed by counting the number of rational points in the intersection of $X$ with each such plane. Let $\Pi$ be one of the planes of bounded height as above. We have to consider some cases. If $\Pi\subseteq X$, suppose without loss of generality that $\Pi$ is the plane $x_{3}=0$. This inclusion allows us to conclude that $F(x_{0},x_{1},x_{2},0)$ is the zero polynomial, hence $x_{3}$ divides $F$, which is then reducible, a contradiction since $X$ is smooth. So this cannot occur. We conclude that $X\cap\Pi$ is a plane curve of degree $d$.
Let $C$ be an irreducible degree $e\leq d$ component of this curve, and let $\tilde{C}$ be its normalization. We can compute the genus of $\tilde{C}$ as being
$g(\tilde{C})=\frac{(e-1)(e-2)}{2}-\sum_{p}\frac{m_{p}(m_{p}-1)}{2},$ (2)
where the sum is taken over all singular points of $C$, including infinitely near singular points, and $m_{P}$ denotes the multiplicity of $C$ at $p$ (see, for example, [7, Example 3.9.2]). In particular, as $g(\tilde{C})\geq 0$, we see that $C$ has a finite number of singular points, and this number is bounded in terms of $d$. Since $\tilde{C}$ coincides with $C$ outside the finite set of singular points, we see that counting the rational points of $C(K)$ and $\tilde{C}(K)$ is the same up to $O_{d}(1)$.
If $g=g(\tilde{C})\geq 2$, we can apply the Uniform Faltings’s Theorem (Lemma 1.1) to $\tilde{C}$ to find
$\tilde{C}(K)\leq c(g)^{1+\rho(\tilde{C})}$.
Let $\varepsilon>0$. We will show that, assuming the Rank Hypothesis, we have
$\tilde{C}(K)\ll_{K,d,\varepsilon}B^{\varepsilon}.$
We start by observing that, since $g\leq\frac{(d-1)(d-2)}{2}$, any dependence on $g$ can be transformed into a dependence on $d$. We will use this fact several times in what follows.
The curve $\tilde{C}$ is obtained by normalizing an irreducible component of the intersection $X\cap\Pi$, where $H_{K}(\Pi)\ll_{K}B^{1/3}$. This means that we have an equation for $\Pi$ of the form
$a_{0}x_{0}-a_{1}x_{1}-a_{2}x_{2}-a_{3}x_{3}=0,$
where $H_{K}([a_{0}\colon a_{1}\colon a_{2}\colon a_{3}])\ll_{K}B^{1/3}$. Supposing without loss of generality $a_{0}\neq 0$, the intersection $X\cap\Pi$ is defined by the equation:
$F_{\Pi}(x_{1},x_{2},x_{3})\coloneqq F(a_{1}x_{1}+a_{2}x_{2}+a_{3}x_{3},a_{0}x_{1},a_{0}x_{2},a_{0}x_{3})=0.$
Write
$F(x_{0},x_{1},x_{2},x_{3})$ $=$ $\sum_{\begin{subarray}{c}e\in\operatorname{\mathbb{N}}^{4},\\ \lvert e\rvert=d\end{subarray}}b_{e}x_{0}^{e_{0}}x_{1}^{e_{1}}x_{2}^{e_{2}}x_{3}^{e_{3}},\text{ and}$
$F_{\Pi}(x_{1},x_{2},x_{3})$ $=$ $\sum_{\begin{subarray}{c}f\in\operatorname{\mathbb{N}}^{3},\\ \lvert f\rvert=d\end{subarray}}c_{f}x_{1}^{f_{1}}x_{2}^{f_{2}}x_{3}^{f_{3}}.$
Expanding the expression for $F_{\Pi}$, we obtain:
$F_{\Pi}(x_{1},x_{2},x_{3})$ $=$ $\sum_{\begin{subarray}{c}e\in\operatorname{\mathbb{N}}^{4},\\ \lvert e\rvert=d\end{subarray}}b_{e}(a_{1}x_{1}+a_{2}x_{2}+a_{3}x_{3})^{e_{0}}x_{1}^{e_{1}}x_{2}^{e_{2}}x_{3}^{e_{3}}$
$=$ $\sum_{\begin{subarray}{c}e\in\operatorname{\mathbb{N}}^{4},\\ \lvert e\rvert=d\end{subarray}}b_{e}\sum_{\begin{subarray}{c}j\in\operatorname{\mathbb{N}}^{3},\\ \lvert j\rvert=e_{0}\end{subarray}}\binom{e_{0}}{j_{1},j_{2},j_{3}}(a_{1}x_{1})^{j_{1}}(a_{2}x_{2})^{j_{2}}(a_{3}x_{3})^{j_{2}}(a_{0}x_{1})^{e_{1}}(a_{0}x_{2})^{e_{2}}(a_{0}x_{3})^{e_{3}}$
$=$ $\sum_{\begin{subarray}{c}e\in\operatorname{\mathbb{N}}^{4},\\ \lvert e\rvert=d\end{subarray}}b_{e}\sum_{\begin{subarray}{c}j\in\operatorname{\mathbb{N}}^{3},\\ \lvert j\rvert=e_{0}\end{subarray}}\binom{e_{0}}{j_{1},j_{2},j_{3}}a_{0}^{e_{1}+e_{2}+e_{3}}a_{1}^{j_{1}}a_{2}^{j_{2}}a_{3}^{j_{3}}x_{1}^{e_{1}+j_{1}}x_{2}^{e_{2}+j_{2}}x_{3}^{e_{3}+j_{3}}$
$=$ $\sum_{\begin{subarray}{c}f\in\operatorname{\mathbb{N}}^{3},\\ \lvert f\rvert=d\end{subarray}}b_{(j_{1}+j_{2}+j_{3},f_{1}-j_{1},f_{2}-j_{2},f_{3}-j_{3})}\sum_{\begin{subarray}{c}j\in\operatorname{\mathbb{N}}^{3},\\ \lvert j\rvert\leq d\end{subarray}}\binom{j_{1}+j_{2}+j_{3}}{j_{1},j_{2},j_{3}}a_{0}^{d-j_{1}-j_{2}-j_{3}}a_{1}^{j_{1}}a_{2}^{j_{2}}a_{3}^{j_{3}}x_{1}^{f_{1}}x_{2}^{f_{2}}x_{3}^{f_{3}}.$
This shows that, for each $f\in\operatorname{\mathbb{N}}^{4}$, $\lvert f\rvert=d$, we have:
$c_{f}=b_{(j_{1}+j_{2}+j_{3},f_{1}-j_{1},f_{2}-j_{2},f_{3}-j_{3})}\sum_{\begin{subarray}{c}j\in\operatorname{\mathbb{N}}^{3},\\ \lvert j\rvert\leq d\end{subarray}}\binom{j_{1}+j_{2}+j_{3}}{j_{1},j_{2},j_{3}}a_{0}^{d-j_{1}-j_{2}-j_{3}}a_{1}^{j_{1}}a_{2}^{j_{2}}a_{3}^{j_{3}}.$
Therefore, for any place $v$ of $K$:
$\max_{\begin{subarray}{c}f\in\operatorname{\mathbb{N}}^{3},\\ \lvert f\rvert=d\end{subarray}}\lVert c_{f}\rVert_{v}\ll_{d}\max_{\begin{subarray}{c}e\in\operatorname{\mathbb{N}}^{4},\\ \lvert e\rvert=d\end{subarray}}\lVert b_{e}\rVert_{v}\cdot\big(\max_{0\leq i\leq 3}\lVert a_{i}\rVert_{v}\big)^{d}.$
Furthermore, if $v$ is a non-archimedean place, the explicit constant in the inequality above can be taken to be $1$. Multiplying the equation above for all $v$, we obtain:
$H_{K}(F_{\Pi})\ll_{d}H_{K}(F)\cdot H_{K}([a_{0}\colon a_{1}\colon a_{2}\colon a_{3}])^{d}\ll_{K}H_{K}(F)\cdot B^{d/3}.$
Taking the logarithm on both sides, we obtain an inequality of logarithmic heights:
$h_{K}(F_{\Pi})\ll_{K,d}h_{K}(F)+\log B+1.$
Hence, denoting by h(C~)h($\tilde${C}) the naive height of the curve $\tilde{C}$ (that is, the minimum logarithmic height of equations defining $\tilde{C}$), we have:
$h(\tilde{C})\leq h_{K}(F_{\Pi})\ll_{K,d}h_{K}(F)+\log B+1.$
Let $h_{\operatorname{Falt}}(J)$ be the Faltings height of $J$. We want to compare $h_{\operatorname{Falt}}(J)$ with h(C~)h($\tilde${C}). For this, we will also use $h_{\Theta}(J)$, the theta height of $J$ relative to the sixteenth power of a symmetric translation of the theta divisor of $J$. By [16, Corollary 1.3]:
$\lvert h_{\Theta}(J)-\frac{1}{2}\cdot h_{\operatorname{Falt}}(J)\rvert\ll_{g}\log(2+\max\{h_{\Theta}(J),1\}).$
This gives us an inequality:
$h_{\operatorname{Falt}}(J)\ll_{d}\max\{h_{\Theta}(J),1\}$
Using [17, Théorème 1.3 and Proposition 1.1] and using that $g$ is bounded in function of $d$, we obtain:
$h_{\Theta}(J)\ll_{d}h(\tilde{C})+1.$
Combining the two inequalities above and using the inequality obtained for the naive height gives:
$h_{\operatorname{Falt}}(J)\ll_{d}h(\tilde{C})+1\ll_{K,d}h_{K}(F)+\log B+1.$
It is well-known that $\log N_{J}\ll h_{\operatorname{Falt}}(J)$ (one can deduce this inequality for Abelian varieties with semistable reduction over $K$ by using [4, Theorem A], and from this deduce the general case by using [9, Lemma 3.4]). Therefore, by the Rank Hypothesis:
$\rho(\tilde{C})=r_{J}=o(\log N_{J})=o(h_{\operatorname{Falt}}(J))=o_{K,d}(h_{K}(F)+\log B).$
Hence, there exists a constant $B_{0}(\varepsilon)>0$ such that, for all $B\geq B_{0}(\varepsilon)$,
$1+\rho(\tilde{C})\ll_{K,d}\varepsilon(h_{K}(F)+\log B).$
Notice that, since we are assuming $H_{K}(F)\ll_{K}B^{d\binom{d+3}{3}}$, we have $h_{K}(F)\ll_{K,d}\log B$. Hence, for $B\leq B_{0}(\varepsilon)$, we have $h_{\operatorname{Falt}}(J)\ll_{K,d}\log B_{0}(\varepsilon)$. Since the number of Abelian varieties of bounded height is finite, there exist constants $c_{1}(K,d),c_{2}(K,d,\varepsilon)>0$ such that
$1+\rho(\tilde{C})\ll_{K,d}c_{1}(K,d)\varepsilon(h_{K}(F)+\log B)+c_{2}(K,d,\varepsilon)$
for all $\tilde{C}$. Thus, by the Uniform Faltings’s Theorem:
$\#\tilde{C}(K)\leq c(g)^{1+\rho(\tilde{C})} \displaystyle\leq$ $c(g)^{c_{1}(K,d)\varepsilon(h_{K}(F)+\log B)+c_{2}(K,d,\varepsilon)}$
$=$ $c(g)^{c_{2}(K,d,\varepsilon)}\cdot c(g)^{c_{1}(K,d)\varepsilon(\log(H_{K}(F)B))}$
$\ll_{K,d,\varepsilon}$ $(H_{K}(F)B)^{\varepsilon c_{1}(K,d)\log c(g)}$
$\ll_{K,d,\varepsilon}$ $B^{\big(d\binom{d+3}{3}+1\big)\varepsilon c_{1}(K,d)\log c(g)}.$
Adjusting the constant $B_{0}(\varepsilon)$ if needed, we obtain
$\tilde{C}(K)\ll_{K,d,\varepsilon}B^{\varepsilon},$
as we wanted.
If $g(\tilde{C})=1$, then either $\tilde{C}(K)=\emptyset$ or $\tilde{C}$ is an elliptic curve. In the first case, there is nothing to do. Suppose that we are in the second case. We will use [6, Proposition 1.4], which we restate here:
Lemma 2.2.
Assume that the Rank Hypothesis holds. Let $E\subseteq\mathbb{P}^{n}$ be a smooth curve of genus $1$. Then, for any $\varepsilon>0$, we have $N_{E}(B)\ll_{K,d,m,\varepsilon}B^{\varepsilon}$.
If $C$ has no smooth points of height smaller than or equal to $B$, then $N_{C}(B)\ll_{d}1$, since $C$ has $O_{d}(1)$ singular points. Otherwise, fix a smooth point $x_{0}\in C$ with $H_{K}(x_{0})\leq B$. Then, by applying Riemann-Roch to divisors of the form $m(x_{0})$, we can find an embedding $\tilde{C}\xhookrightarrow{}\mathbb{P}^{2}$ with respect to which
$h(\tilde{C})\ll h(C)+\log H_{K}(x_{0})\ll\log B,$
where h(C~)h($\tilde${C}) and $h(C)$ denote the naive heights of $\tilde{C}$ and $C$ with respect to the embeddings $\tilde{C}\xhookrightarrow{}\mathbb{P}^{2}$ and $C\xhookrightarrow{}\mathbb{P}^{3}$, respectively.
We can take $m=2$ in the above lemma to get $N_{\tilde{C}}(B)\ll_{K,d,\varepsilon}B^{\varepsilon}$, with the height induced by the map $\tilde{C}\xhookrightarrow{}\mathbb{P}^{2}$. The above comparison of heights allows us to conclude that $N_{C}(B)=O_{K,d,\varepsilon}(B^{\varepsilon})$ for any $\varepsilon>0$. Therefore, the total contribution of the points that come from these curves is $O_{K,d,\varepsilon}(B^{4/3+\varepsilon})$.
All that is left is to bound the number of rational points in curves of genus 0 that lie inside $X$. If $C\subseteq X$ is a curve of genus 0, then either $C(K)=\emptyset$ or $C$ is a rational curve. In the first case, there is nothing to do. In the second case, let $e$ be the degree of $C$. Then $1\leq e\leq d$. If $e=1$, then $C$ is a line, and by definition it does not contribute to the calculation of $N_{X^{\prime}}(B)$. Suppose $e\geq 2$. By [15, Theorem 1.8],
$N_{C}(B)\ll_{K,d,\varepsilon}B^{2/e+\varepsilon}\leq B^{1+\varepsilon}.$
Since the contribution of a single rational curve of degree $e\geq 2$ is less than $O_{K,d,\varepsilon}(B^{4/3+\varepsilon})$, and since we have $O_{d}(1)$ possibilities for $e$, we conclude the proof of Theorem 1.3 if we show the following:
Lemma 2.3.
Let $X\subseteq\mathbb{P}^{3}$ be a smooth surface of degree $d\geq 4$. Fix a positive integer $e$. Then $X$ contains finitely many rational plane curves of degree $e$. Furthermore, the number of such curves is $O_{d}(1)$.
Proof.
Assume by contradiction that $X$ contains infinitely many rational plane curves of degree $e$. A plane curve of degree $e$ has arithmetic genus $\frac{(e-1)(e-2)}{2}$, hence the Hilbert polynomial of any such curve is given by
$P(m)=m\cdot e+1-\frac{(e-1)(e-2)}{2}.$
Consider the Hilbert Scheme $\operatorname{Hilb}_{X}^{P}$ that parameterizes the subvarieties of $X$ with Hilbert Polynomial $P$. It is a projective scheme (see [11, Section I.1]). Notice that the irreducible plane curves of degree $e$ inside $X$ form an open set $\operatorname{Irr}_{X}^{e}\subseteq\operatorname{Hilb}_{X}^{P}$. Let $\operatorname{Rat}_{X}^{e}\subseteq\operatorname{Irr}_{X}^{e}$ be the set of rational plane curves of degree $e$ inside $X$. By (2), being a rational curve is a closed condition in $\operatorname{Irr}_{X}^{e}$. Hence $\operatorname{Rat}_{X}^{e}$ is a closed subvariety of $\operatorname{Irr}_{X}^{e}$, and therefore it is a quasi-projective scheme. In particular, it has finitely many irreducible components.
By our hypothesis, $\operatorname{Rat}_{X}^{e}$ has infinitely many points, hence there is an irreducible component $T\subseteq\operatorname{Rat}_{X}^{e}$ which contains infinitely many points. For any $t\in T$, let $C_{t}$ be the curve associated with $t$. Let $U\subseteq X\times T$ be the universal family over $T$, that is,
$U=\{(x,t)\colon x\in C_{t}\}.$
Consider the projection in the first coordinate $\pi\colon U\to X$. Suppose that this map is not dominant. Then its image $\bigcup_{t\in T}C_{t}$ must lie inside a proper subvariety of $X$. As $X$ is a surface, this means $\bigcup_{t\in T}C_{t}$ is contained in a curve. But this is impossible, since such a curve would have infinitely many irreducible components.
We conclude that $\pi\colon U\to X$ is dominant, that is, $X$ is covered by a family of rational curves. This shows that $X$ is uniruled, and thus it has Kodaira dimension -$\infty$.
However, as $X$ is a smooth surface of degree $d\geq 4$, it is either a K3 surface (if $d=4$) or a surface of general type (if $d\geq 5$). In the first case, $X$ has Kodaira dimension 0, and in the second case it has Kodaira dimension $2$. This contradiction proves the finiteness of $\operatorname{Rat}_{X}^{e}$.
To conclude the proof of the Lemma, it suffices to show that the degree of $\operatorname{Rat}_{X}^{e}$ is $O_{d}(1)$. Define $\operatorname{Rat}_{\mathbb{P}^{3}}^{e}$ as the set of rational plane curves of degree $e$ inside $\mathbb{P}^{3}$. By the same argument as above, $\operatorname{Rat}_{\mathbb{P}^{3}}^{e}$ is a quasi-projective scheme. Clearly, $\operatorname{Rat}_{X}^{e}=\operatorname{Rat}_{\mathbb{P}^{3}}^{e}\cap\operatorname{Hilb}_{X}^{P}$, hence by Bézout’s Theorem it suffices to show that the degree of $\operatorname{Hilb}_{X}^{P}$ is $O_{d}(1)$.
Let $N=\binom{d+3}{3}-1$, and $\mathbb{P}^{N}$ be the moduli space of surfaces of degree $d$ contained in $\mathbb{P}^{3}$. Consider the incidence variety
$I=\{(Y,C)\in\mathbb{P}^{N}\times\operatorname{Hilb}_{\mathbb{P}^{3}}^{P}\colon C\subseteq Y\}.$
We have a projection $p\colon I\to\mathbb{P}^{N}$, with respect to which $p^{-1}(X)\cong\operatorname{Hilb}_{X}^{P}$. As the degree of a variety is stable by algebraic deformation, we see that $\operatorname{Hilb}_{X}^{P}$ has degree $O_{d}(1)$, as we wanted. ∎
This concludes the proof of Theorem 1.3.
3. Proof of Theorem 1.4
We can adapt our approach to bound the number of rational points of bounded height on surfaces lying inside a larger projective space. Let $X\subseteq\mathbb{P}^{n}$ be a non-degenerate smooth surface of degree $d$ over $K$. The main difference here is that the ideal of $X$ is generated by several forms. Instead of Lemma 2.1, we use [2, Lemma 5], and compare $\lVert\cdot\rVert$ with $H_{K}(\cdot)$ as before by using the sets of representatives $Z_{m}$.
By [19, Theorem 2], the set of points of $X(K)$ of height smaller than $B$ can be covered by hyperplanes of height $O_{K,n}(B^{1/n})$, and by [19, Theorem 3] there are $O_{K,n}(B^{\frac{n+1}{n}})$ such planes.
Since $X$ is non-degenerate, for each such hyperplane $\Pi$ the intersection $X\cap\Pi$ is a curve $C$ of degree $d$. By Castelnuovo’s bound, $p_{a}(C)\ll_{n,d}1$, where $p_{a}(C)$ denotes the arithmetic genus of $C$. By [7, Example 3.9.2],
$g(C)=p_{a}(C)-\sum_{p}\frac{m_{p}(m_{p}-1)}{2},$
This shows that the number of singularities of $C$ is $O_{n,d}(1)$. The proof now follows identically to the proof of Theorem 1.3, and it suffices to show the following generalization of Lemma 2.3:
Lemma 3.1.
Let $X\subseteq\mathbb{P}^{n}$ be a non-uniruled surface of degree $d$. Fix a positive integer $e$. Then $X$ contains finitely many rational curves of degree $e$. Furthermore, the number of such curves is $O_{n,d}(1)$.
Proof.
It suffices to apply the argument in the proof of Lemma 2.3 to any possible Hilbert polynomial of a rational curve of degree $e$ contained in $X$. By Castelnuovo’s bound, there are finitely many such polynomials. We conclude, as in the proof of Lemma 2.3, that if $X$ contained infinitely many rational curves of degree $e$, then $X$ would be uniruled, proving the lemma. ∎